### v1

# v-Hub: A Visual-Centric Humor Understanding Benchmark for Video LLMs

Abstract
AI models capable of comprehending humor hold real-world promise—for example, enhancing engagement in human-machine interactions. To gauge and diagnose the capacity of multimodal large language models (MLLMs) for humor understanding, we introduce v-Hub, a novel visual-centric video humor understanding benchmark. v-Hub comprises a curated collection of minimally verbal short videos, sourced from classic silent films and online resources, and reflecting real-world scenarios where humor can be appreciated purely through visual cues. Each video clip is paired with rich annotations, including captions, descriptions, and explanations, supporting evaluation tasks like caption matching and humor explanation. To broaden its applicability, we further construct an open-ended video QA task, making it readily integrable into existing video understanding benchmarks. We evaluate a diverse set of MLLMs, from specialized Video-LLMs to versatile OmniLLMs that can process audio, covering both open-source and proprietary domains. The experimental results expose the difficulties MLLMs face in comprehending humor from visual cues alone. For example, all models exhibit a marked performance drop on caption matching when moving from text-based to video-based evaluation (without audio). Our findings also demonstrate that incorporating audio helps with video humor understanding, highlighting the informativeness of sound and the promise of integrating richer modalities for complex video understanding tasks.
: Equal Contribution. $\dagger$: Project Lead. ✉: Corresponding Author. Contact: shi_zpeng@sjtu.edu.cn,lihengli@stu.pku.edu.cn,yannzhao.ed@gmail.com.
(a) Visuals. A man placed a battery on the conveyor belt, but it rolled against the belt’s motion, forcing the cashier into an endless wait. For those who know the physics of a rolling cylinder on a moving conveyor, the scene feels even more clever.
(b) Visuals+Text. The video shows an animal rescue, with a cow dangling beneath a helicopter, appearing to swirl midair. The scene seems routine at first, but the added text ‘milkshakes’ cleverly parallels the moment, making it unexpectedly witty.
(c) Visuals+Audio. As the man flips through the pages, cartoon characters gradually appear, accompanied by a distinct melody. First, the dancer’s rhythm and the suona player’s piercing tune, then the cymbal player’s resonant clash, together creating an evolving effect.
(d) Visuals+Audio+Text. A guy messaged his friend that he was making a birthday cake for them. After it was baked and sliced, the inside mimicked their chat bubble layout. The whole scene was made even merrier by the Happy Birthday melody.
Figure 1: Examples of visual-centric humor understanding, where ‘audio’ and ‘text’ refer to environmental sound (cf. human speech) and visual text, respectively.
1 Introduction
Humor enriches our daily lives and appears in many forms, from jokes and cartoons to comedies and viral videos. AI models capable of understanding humor hold promise for engaging with humans empathetically (Hampes, 2001; 2010), but perceiving and comprehending humor can be challenging even to humans due to the heavy reliance on nontrivial reasoning, social and cultural contexts, etc (see Figure 1). This, on the other hand, makes humor understanding a promising testbed to evaluate how well state-of-the-art AI models understand humor. Indeed, there has been a line of research centering around gauging the capability of pre-trained large language models (LLMs) for humor understanding (Hessel et al., 2022; Hyun et al., 2023; Ko et al., 2023), but parallel work on multimodal LLMs is still lacking, though they are more naturally suited for understanding multimodal humor.
In this work, we address this gap by investigating humor understanding with multimodal LLMs (MLLMs), focusing specifically on MMLMs that are capable of processing video. We choose video as the primary medium of humor, since it captures nuanced variations and diverse styles, presenting a unique challenge for MLLMs. For example, perceiving the humor in Figure1(d) requires recognizing visual text and the layout of chat bubbles and understanding their temporal and semantic correspondences with the cut surface of the cake slice. While there have been a few benchmarks containing humorous videos (see Table 1), all of them were designed exclusively for the evaluation of LLMs (Ko et al., 2023; Hyun et al., 2023). [^1] They translated videos into language descriptions and performed verbal humor evaluation with LLMs. Moreover, they are limited in that each humor either is dominated solely by spoken language (Hyun et al., 2023) or can be understood only when both the video and linguistic cues are present (Ko et al., 2023), ignoring the fact that humans can understand humor from visual cues alone, exemplified by Charlie Chaplin’s silent comedies.
To address this limitation, we curate a set of visual-centric humorous videos from two complementary sources: Charlie Chaplin’s silent films and user-generated short funny videos. Silent film humor is conveyed through visual cues, but is thematically and culturally constrained due to the scripted performance. To increase diversity, we incorporate user-generated funny short videos from various occasions and cultural backgrounds. We rigorously filtered the videos to retain only those where the humor is primarily visual. Our final dataset consists of videos where humor is derived predominantly from the visual modality (e.g., 99% of all videos), making it visual-centric and more suitable for diagnosing the visual reasoning ability of MLLMs.
To assess how well MLLMs understand humor in video, we create a visual-centric humor understanding benchmark (v-Hub), which consists of three distinct tasks. (1) First, the Caption Matching task challenges MLLMs to align video captions with the corresponding videos. Apart from testing surface-level matching, the task is carefully designed to require an appreciation of nuanced, extended humor. (2) Second, the Humor Explanation task evaluates whether MLLMs can extract humor elements and provide accurate rationales. (3) Finally, the Open-ended QA task evaluate the MLLMs’ fundamental understanding of videos from humor genre across temporal, descriptive, and causal dimensions, broadening the applicability of v-Hub. Together, these tasks provide a comprehensive framework to benchmark MLLMs in visual-centric humor understanding.
We evaluate representative MLLMs from both open- and closed-source domains. Depending on the input modalities, we consider the following three task settings. (1) The Text-Only setting assumes human-level interpretation of video contents and provides detailed human-written descriptions. (2) The Video-Only setting offers only videos (without audio) to assess the ability of MLLMs to derive humor solely from visual cues. (3) We further propose a novel Video+Audio setting that combines visual and auditory signals to determine whether sound cues—such as background music and sound effects—help MLLMs (aka. OmniLLMs) better understand humor. [^2] In this work, audio primarily refers to environmental sound rather than human speech (see Section 2.2).
We empirically find that MLLMs generally perform better with text-only inputs than with video-only inputs (see Table 2). For example, Qwen2.5-VL-72B drops in accuracy from 0.719 to 0.673 on Caption Matching, and Gemini-2.5-Flash from 0.611 to 0.583, under the video-only setting, indicating their struggles in capturing subtle visual cues for humor understanding. Adding audio yields slight improvements across most OmniLLMs. For instance, MiniCPM-2.6-o improves from 0.364 to 0.404 in accuracy on Caption Matching, confirming the effectiveness of the audio modality, though it still lags behind the text-only setting. Overall, our v-Hub presents a new challenge and contributes to a comprehensive evaluation of MLLMs. It exposes their weakness in visual-centric humor understanding, stresses the need for enhancing their visual reasoning capabilities, and highlight the promise of integrating additional modalities like sound for video understanding.
Table 1: Comparison between v-Hub and prior humor video datasets.
Dataset Type Visual-Centric Tasks
Humor Explanation Caption Matching Open-ended QA
NYCC (Hessel et al., 2022) Cartoon [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MUStARD (Castro et al., 2019) Sitcom [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
WITS (Kumar et al., 2022) Sitcom [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
UR-FUNNY (Hasan et al., 2019) TED Talks [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
SMILE (Hyun et al., 2023) Sitcom, TED Talks [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
ExFunTube (Ko et al., 2023) Short videos [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
v-Hub (ours) Short videos, Silent films [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
2 Curating Visual-Centric Humorous Videos
2.1 Humor Video Sources
Our goal is to collect humorous videos that are visual-centric and illustrate diverse humor. A straightforward approach is to collect humorous clips from silent comedies that are entirely devoid of speech. Though silent films may contain recorded music, sound effect, and few captions, which may contribute to the expression of humor, the humor primarily arises from the visual modality. A major issue with silent film clips is that they have rather narrow themes and employ limited storytelling techniques. To enhance the diversity of humor in our dataset, we further incorporate user-generated short funny videos from the Internet. Specifically, we selected videos from an X account (@humansnocontext) that frequently shares humorous clips with minimal reliance on speech or text-based context. Thus, our dataset comprises humorous videos from two different domains that complement each other (see Figure 2):
  • Charlie Chaplin’s Silent Films: We reviewed Charlie Chaplin’s classic silent films from 1914 to 1938 and collected 729 funny clips. Each humor is ensured to be self-contained, without relying on additional video contexts. Figure 2 shows an example in this domain.
  • User-Generated Funny Videos: We reviewed the X user @humansnocontext’s tweets posted between March 28, 2023 and October 12, 2024 and collected 18080 short funny videos.
2.2 Preprocessing and Filtering
We preprocess and filter the initially collected videos according to duration, appropriateness, and speech reliance, sequentially. (1) Duration: We retain videos ranging from 5 to 60 seconds long. Short clips under 5 seconds generally fail to convey meaningful humor, while clips exceeding 1 minute often rely on dialogue. For silent films, we segment long scenes to isolate individual humorous moments, ensuring that each segment captures the full humor, without becoming too long for generation tasks. (2) Appropriateness: To ensure that the contents of our videos are appropriate, we adhered to the safety objectives outlined in Thoppilan et al. (2022) and excluded videos that violated the established criteria (see details in Appendix B.1). (3) Speech reliance: We minimize reliance on speech. Since there is little to no speech in Charlie Chaplin’s silent films, we primarily focused on user-generated funny videos and employed both manual and automatic approaches to filter out speech-heavy videos (see details in Appendix B.1).
Figure 2: Data Curation Pipeline. To collect visual-centric humorous videos, the pipeline consists of two main stages: (a) Humorous video collection, where annotators identify timestamps of self-contained humorous clips for silent films and verify humor presence in short videos (see Section 2.1). (b) Filtering and annotation, where only visual-dominant humor is retained and annotated (see Section 2.3). The annotation is further used for task construction (see Section 3).
2.3 Annotation
We recruited eight annotators based on the following criteria: (1) sufficient English proficiency to understand video content, (2) broad cultural knowledge to interpret humor arising from various contexts, and (3) strong observational skills assessed through a qualification test (see Appendix B.2). To ensure consistency, we provided detailed guidelines for each annotation task and created a reference manual for on-demand use. Each video underwent three rounds of annotation to guarantee correctness and thoroughness. We conducted the following primary annotation tasks (see Figure 3 for an example annotation):
  • Humor Evaluation: Annotators independently evaluated whether the video was humorous.
  • Captioning: Each annotator was asked to write two types of captions for each video, without seeing existing annotations, including captions and descriptions, from other annotators, thus ensuring an independent and unbiased judgment.
    • 1.
      Descriptive Captions directly describe or highlight the humor present in the video content from the original publisher’s perspective.
    • 2.
      Creative Captions extend beyond the video’s original humor by adding imaginative or novel elements (see the added visual caption in Figure 1(b)).
    The dual-caption annotation supports a comprehensive assessment of humor in video from both comprehension and generation perspectives (see Section 3).
  • Video Description: Annotators were instructed to describe the events in each video without making inferences, focusing only on observable objects, actions, and expressions. After the first annotator completes the video description, subsequent annotators review and refine these descriptions for correctness and completeness.
  • Video Labeling: Annotators labeled the key humor sources (e.g., human actions, objects, visual effects, or sound cues) in each video and noted whether any visual text was present. If an element appeared, but did not contribute to humor, it was not selected.
  • Humor Explanation: Three annotators sequentially create and refine humor explanations by adding missing details, guaranteeing comprehensive coverage of the labeled humor sources through an iterative refinement process.
2.4 Data Analysis
After all filtering processes, we were left with 960 videos, including 267 silent humor clips from Charlie Chaplin and 693 user-generated short funny videos from the Internet. The total duration of the videos is 4h, and the average duration is around 15s. All of them rely on the visual modality to express humor. We identify two key modalities that dominate the delivery of humor: visuals and audio. Apart from 600 videos (63%) conveying humor primarily via pure visual cues (denoted by ‘Visual’), 92 videos (9%) contain additional linguistic cues in visual form—such as embedded captions and subtitles (denoted by ‘Visual+Text’)—that extend humor, 214 video humor (22%) is enhanced by additional sound that covers non-speech auditory elements, such as background music, sound effects, and character vocalizations (denoted by ‘Visual+Audio’), and 46 videos (5%) convey humor through visuals, sound, and visual text (denoted by ‘Visual+Audio+Text’). The video distribution over the four groups is illustrate in Figure 8 in Appendix C, and more analysis can be found there.
Figure 3: Example annotation of a short video that conveys humor through visuals, visual text, and background sound. Knowing the Happy Birthday melody makes the video merrier (see Section 2.3).
3 v-Hub: a visual-centric humor understanding benchmark
3.1 Evaluation tasks
To comprehensively evaluate the capability of MLLMs in humor understanding, we propose three tasks that reflect different aspects of humor reasoning: Caption Matching, Humor Explanation, and Open-ended QA:
  • Caption Matching. In this discriminative task, models must correctly associate videos with their corresponding captions. Unlike ordinary caption matching tasks, our design challenges MLLMs to go beyond surface-level matching and assess their ability to understand video humor that is pronounced by creative captions from a generation perspective. For each video with a creative caption, we randomly sample four descriptive captions from other videos as the distractors.
  • Humor Explanation. In this generative task, models must identify humor points within each video, provide coherent explanations, and reference relevant visual or auditory cues.
  • Open-ended QA. To further assess the fundamental understanding of video content, we generate a set of open-ended question-answer pairs for each video (see details in Appendix D.1). These questions—automatically generated by GPT-4o (Hurst et al., 2024) and manually verified—encompass temporal, descriptive, and causal aspects (Xiao et al., 2021). [^3] There are 62, 675, and 223 QA pairs for temporal, descriptive, and causal questions, respectively. This extends the benchmark beyond humor-specific reasoning, providing a broader assessment of video reasoning skills.
3.2 Evaluation Methods
We employ different evaluation strategies depending on the task type:
  • Accuracy. For the caption matching task, we measure accuracy to determine whether the model correctly identifies the most appropriate response.
  • Quality of Open-ended Responses. For humor explanation and open-ended QA tasks, we adopt both automatic and human evaluation approaches:
    • Semantic Similarity. We compute similarity scores between model-generated answers and human-provided answers using BERTScore (Zhang* et al., 2020), which captures fine-grained semantic similarity beyond simple word overlap. It provides three metrics: recall, precision, and F1 score (see Appendix D.2 for details). In addition, In addition, we employ SentBERT (Reimers & Gurevych, 2019) to assess sentence-level semantic coherence, as well as METEOR (Banerjee & Lavie, 2005), which provides a more nuanced assessment of semantic adequacy and fluency.
    • Human Evaluation. We randomly sample a subset of model-generated explanations and compare them with human-written explanations. The evaluators rate the explanations based on accuracy and logicality, providing insight into the gap between human and MLLMs’ explanations. Results are presented in Table 9 in Appendix E.
4 Experiments
4.1 Experimental Setup
MLLMs.
We consider both proprietary and public MLLMs like Gemini-2.5-Flash (Team et al., 2025) and Qwen2.5-VL (Bai et al., 2025). OmniLLMs such as Video-SALMONN-2 (Tang et al., 2025) and Qwen-2.5-Omni (Xu et al., 2025), which can process audio, are also included (an overview of all evaluated MLLMs is presented in Table 6).
Evaluation settings.
To understand the roles of different modalities in video humor understanding, we consider the following three settings: Text-Only, Video-Only, and Video+Audio, which means models are tested with text, video (w/ audio), and video-audio inputs, respectively.
  • Text-Only. In this setting, models receive detailed human-written video descriptions; no visual or audio information is available to the models. Thus, it evaluates the language reasoning ability of MLLMs in isolation.
  • Video-Only. Models are provided with only raw video frames, without audio. This setting assesses their intrinsic visual comprehension capabilities. Depending on the presence of visual text, we further divide results into two groups: ‘w/ visual text’ and ‘w/o visual text’.
  • Video+Audio. Models receive both video frames and audio signals, allowing us to examine whether the inclusion of auditory information improves humor understanding. Depending on the contribution of audio to humor, we further divide results into two groups: ‘w/ humor audio’ and ‘w/o humor video’.
4.2 Main Results
Based on the results in Table˜2, we analyze the humor competence of MLLMs along three dimensions: video humor discovery, understanding, and subtle humor inference. Our results reveal several shortcomings of MMLMs: they (i) struggle to identify humorous elements when explicit cues are absent, (ii) inadequately fuse information across modalities for understanding, and (iii) show limited capacity for inferring subtle humor.
Limited ability in humor discovery.
Across settings, models consistently perform better on Open-ended QA than on Humor Explanation. This performance disparity reveals that they are limited in perceiving humor. For example, in the Text-Only setting, Qwen-2.5-VL-72B, whose score drops from 0.792 in QA to 0.553 in Humor Explanation. These findings suggest that models are more successful when the question itself provides explicit cues that direct attention to a specific humorous element in the scene. By contrast, the Humor Explanation task, which requires models to independently identify and articulate the source of humor without such guidance, poses a greater challenge. This indicates that while MLLMs are often able to reason about humor once it is highlighted for them, they struggle with the more cognitively demanding task of discovering humor directly from contextual cues.
Heavy reliance on linguistic cues for humor understanding.
Comparing text-based and video-based evaluations, we observe marked differences across all three tasks, where the Text-Only setting yields substantially higher scores than the video-based settings, implying that current MLLMs are heavily dependent on linguistic cues for humor understanding. For example, On Open-ended QA, Qwen-2.5-VL-72B achieves a SentBERT score of 0.792 with text input, but it plummets to 0.459 when presented with raw video (w/o audio). While the addition of audio provides a marginal but consistent performance boost, this gain is minimal compared to the contribution of text. This wide performance gap suggests that MLLMs’ cross-modal fusion capabilities are still underdeveloped, leading them to rely predominantly on linguistic cues rather than effectively integrating visual and auditory signals. Thus, future work is well-suited for enhancing MLLMs beyond language understanding.
Incapability for subtle humor inference.
The Caption Matching task goes beyond surface-level linking between literal descriptions and videos; instead, it requires models to find the creative caption that enhances or extends humor in the video. We find that most models exhibit limited performance (e.g., below 0.8), suggesting their incompetence for subtle humor inference. For example, under the most favorable conditions, that is, in the Text-Only setting, the top-performing model, GPT-4o, achieves an accuracy of only 0.767. The difficulty is magnified when models must process raw video data. For example, video-SALMONN-2’s accuracy falls sharply from 0.367 in the Text-Only setting to 0.255 in the Video+Audio condition. This pronounced struggle to connect creative, non-obvious text to original visual humor context reveals a critical weakness in the models’ capacity for the abstract, implicit cross-modal reasoning that is fundamental to comprehending sophisticated humor.
Table 2: Model performance on Humor Explanation, Caption Matching, and Open-ended QA.
Explanation Matching Open-ended QA
MLLMs SentBERT METEOR F1 Score Accuracy SentBERT METEOR F1 Score
Text-Only
Gemini-2.5-flash 0.556 0.248 0.580 0.611 0.748 0.652 0.723
Video-SALMONN-2 0.575 0.242 0.589 0.367 0.602 0.445 0.639
MiniCPM2.6-o 0.558 0.239 0.562 0.518 0.578 0.467 0.543
Qwen-2.5-Omini 0.547 0.232 0.570 0.644 0.740 0.555 0.698
Qwen-2.5-VL-72B 0.553 0.250 0.578 0.719 0.792 0.622 0.738
Intern3.5-VL 0.567 0.255 0.580 0.632 0.721 0.578 0.689
GPT-4o 0.569 0.256 0.581 0.767 0.720 0.666 0.718
Video-Only
Gemini-2.5-flash 0.469 0.200 0.550 0.583 0.434 0.275 0.556
video-SALMONN-2 0.281 0.150 0.504 0.259 0.311 0.160 0.525
MiniCPM2.6-o 0.387 0.164 0.520 0.364 0.323 0.110 0.452
Qwen-2.5-Omini 0.388 0.157 0.521 0.55 0.385 0.104 0.488
Qwen-2.5-VL-72B 0.452 0.188 0.547 0.673 0.459 0.201 0.550
Intern3.5-VL 0.433 0.186 0.543 0.640 0.393 0.230 0.542
GPT-4o 0.478 0.198 0.547 0.667 0.431 0.300 0.556
Video+Audio
Gemini-2.5-flash 0.472 0.200 0.550 0.588 0.428 0.275 0.554
video-SALMONN-2 0.296 0.176 0.506 0.255 0.319 0.180 0.538
MiniCPM2.6-o 0.419 0.176 0.523 0.404 0.348 0.245 0.513
Qwen-2.5-Omini 0.442 0.177 0.531 0.623 0.439 0.164 0.529
4.3 Further Analysis
Table 3: The impacts of audio (i.e., sound) and visual text on video humor understanding.
Sound contributing to humor Sound not contributing to humor
Explanation Matching Open-ended QA Explanation Matching Open-ended QA
Models SentBERT METEOR Accuracy SentBERT METEOR SentBERT METEOR Accuracy SentBERT METEOR
w/ visual text
Gemini-2.5-flash 0.532 0.212 0.630 0.488 0.359 0.509 0.219 0.739 0.476 0.319
video-SALMONN-2 0.292 0.190 0.261 0.319 0.189 0.280 0.176 0.293 0.318 0.189
MiniCPM2.6-o 0.490 0.189 0.348 0.374 0.289 0.474 0.197 0.489 0.380 0.282
Qwen-2.5-Omni 0.512 0.192 0.783 0.525 0.176 0.467 0.190 0.75 0.453 0.166
w/o visual text
Gemini-2.5-flash 0.486 0.202 0.523 0.409 0.265 0.455 0.194 0.215 0.423 0.266
video-SALMONN-2 0.296 0.181 0.243 0.299 0.168 0.298 0.173 0.178 0.325 0.182
MiniCPM2.6-o 0.451 0.178 0.341 0.355 0.271 0.393 0.170 0.215 0.334 0.226
Qwen-2.5-Omni 0.471 0.176 0.551 0.457 0.150 0.422 0.173 0.197 0.422 0.166
To conduct a deeper analysis of model results, we further divide our experimental results based on previously annotated humor modalities and background knowledge essential for delivering humor in video, to analyze how different types of humor affect models’ explanatory capability.
Both audio and visual text help with humor understanding.
As shown in Table 3, MLLMs perform better on videos containing visual text or subtitles than on those without linguistic cues under the Video+Audio setting. For example, Gemini-2.5-Flash attains a SentBERT score of 0.532 and a METEOR score of 0.212 for humor explanation with visual text, compared to 0.486 and 0.202 without visual text. When sound does not contribute to humor, the advantage of visual text becomes even more pronounced: Gemini-2.5-Flash improves from 0.455 to 0.509 in Explanation SentBERT and from 0.215 to 0.739 in Matching Accuracy with visual text. These results indicate that while both audio and visual text help with humor understanding, MLLMs rely more heavily on textual cues, and the presence of visual text can effectively compensate for the absence of informative sound.
Table 4: The impact of background knowledge on video humor understanding.
Explanation Matching Open-ended QA
MLLMs SentBERT METEOR F1 Score Accuracy SentBERT METEOR F1 Score
w/ Background Knowledge
video-SALMONN-2 0.468 0.173 0.563 0.331 0.397 0.195 0.535
MiniCPM-2.6-o 0.518 0.203 0.557 0.445 0.430 0.204 0.520
Qwen-2.5-Omni 0.514 0.197 0.557 0.667 0.496 0.220 0.556
w/o Background Knowledge
video-SALMONN-2 0.287 0.178 0.515 0.261 0.300 0.174 0.528
MiniCPM-2.6-o 0.444 0.181 0.540 0.412 0.351 0.252 0.509
Qwen-2.5-Omni 0.462 0.181 0.545 0.630 0.445 0.158 0.526
Table 5: The impact of video era on video humor understanding.
Explanation Matching Open-ended QA
MLLMs SentBERT METEOR F1 Score Accuracy SentBERT METEOR F1 Score
Last-Century Charlie Chaplin’s Silent Films
Gemini-2.5-flash 0.422 0.188 0.541 0.562 0.386 0.221 0.545
video-SALMONN-2 0.281 0.146 0.509 0.165 0.296 0.154 0.513
MiniCPM2.6-o 0.343 0.150 0.508 0.307 0.314 0.128 0.470
Qwen-2.5-Omni 0.339 0.144 0.510 0.494 0.337 0.119 0.493
Contemporary User-Generated Funny Video
Gemini-2.5-flash 0.487 0.205 0.553 0.592 0.452 0.295 0.560
video-SALMONN-2 0.280 0.151 0.503 0.296 0.315 0.163 0.529
MiniCPM2.6-o 0.404 0.170 0.524 0.385 0.326 0.104 0.446
Qwen-2.5-Omni 0.407 0.162 0.525 0.571 0.404 0.098 0.487
Knowledge-based cues facilitate humor understanding.
We identified 357 videos that require contextual background knowledge and evaluated MLLMs under two settings: with and without the explicit provision of such knowledge. As shown in Table 4, MLLMs consistently achieve higher performance when background knowledge is provided. For instance, Qwen-2.5-Omni attains a SentBERT score of 0.514 and an Explanation F1 of 0.557 with background knowledge, compared to 0.462 and 0.545 without. These findings suggest that while MLLMs implicitly encode certain aspects of cultural context, their comprehension of humor is significantly enhanced by the explicit provision of background knowledge, underscoring the central role of linguistic and knowledge-based cues in complex video humor understanding tasks.
MLLMs have greater difficulty in comprehending humor in historically distant videos.
We analyze the performance of MLLMs under the Video-Only setting across two subsets from distinct eras: last-century Charlie Chaplin’s silent films (CCSF) and contemporary user-generated funny videos (UGFV). As shown in Table 5, MLLMs consistently achieve higher scores on UGFV across all evaluation metrics. For example, Gemini-2.5-flash attains an F1 score of 0.553 for Humor Explanation and 0.560 for Open-ended QA on UGFV videos, compared to 0.541 and 0.545, respectively, on CCSF videos. These findings suggest that MLLMs face greater difficulty in comprehending humor in historically distant videos, highlighting the sensitivity of humor understanding to the temporal and cultural context of videos.
5 Related Work
Video LLMs.
Video LLMs have shown remarkable performance in many traditional video processing tasks such as video captioning (Xu et al., 2016; Agrawal et al., 2019; Plummer et al., 2017), video question answering (Antol et al., 2015; Xiao et al., 2021; Yu et al., 2019; Fu et al., 2025), and grounding (Kazemzadeh et al., 2014; Wu et al., 2022). However, most existing benchmarks primarily target general video understanding tasks, such as MVBench (Li et al., 2024), Video-MME (Fu et al., 2025), PerceptionTest (Patraucean et al., 2023), MLVU (Zhou et al., 2025), and LVBench (Wang et al., 2024b), which mainly assess the recognition of basic visual cues across videos of varying lengths. Others are designed to evaluate specific video understanding capabilities, including temporal grounding (Gao et al., 2017; Lei et al., 2021; Hendricks et al., 2017; Wang et al., 2024c), video object detection (Shang et al., 2019; 2017), and video hallucination (Wang et al., 2024d; Leng et al., 2024). But there remains a pressing need for benchmarks that evaluate higher-level cognitive abilities, such as social intelligence, in order to better measure the gap between human and MLLMs’ performance.
Our work narrows this gap. We expand the evaluation spectrum of video LLMs by introducing a novel humor understanding evaluation framework, formulating a humor generation task, and presenting a first comprehensive evaluation.
Humor Video Understanding.
Humor understanding has been a popular research topic in the area of artificial intelligence and has its roots in cognitive science (Hampes, 2001; 2010),. Early works focus on verbal humor in the form of jokes, sarcasm, etc (Chłopicki, 2005; Petrović & Matthews, 2013; Joshi et al., 2017). As AI models become capable of processing more data modalities like images, videos, and audio, many efforts have been devoted to multimodal humor understanding. For example, Hessel et al. (2022) analyze cartoon images and humorous captions, Desai et al. (2022); Kumar et al. (2022) investigate sarcasm, a special humor type, with image-language data, and Castro et al. (2019); Hasan et al. (2019); Kayatani et al. (2021); Patro et al. (2021); Alnajjar et al. (2022) focus on laughter detection and explanation with videos, including video presentations, sitcoms, and stand-up comedy, but laughter is not the only emotion reaction to humor.
With the development of LLMs, recent works tested how well LLMs understand multimodal humor with image-language and video-language humor (Hessel et al., 2022; Alnajjar et al., 2022), but a parallel evaluation of video LLMs is still missing. While previous works have introduced several video-based humor datasets (Kumar et al., 2022), humor in their videos is either primarily dominated by spoken dialogue or restricted to those that have to be understood relying on both visual and linguistic cues. In contrast, we introduce a visual-centric humor video dataset designed to simulate common scenarios where humans can understand humor purely from visual cues.
Going beyond visual and verbal humor understanding, sound has been found informative of commonsense (Zhao et al., 2022; Zellers et al., 2022), and several studies demonstrate that integrating textual, acoustic and visual characteristics can significantly improve humor detection accuracy (Chandrasekaran et al., 2016; Hasan et al., 2019). Since multimodal LLMs have recently been extended to support audio processing (aka. OmniLLMs), we propose and conduct a first evaluation of MLLMs on video humor understanding that involves sound.
6 Conclusion
We have introduced v-Hub, a visual-centric humor understanding benchmark. v-Hub is designed to assess and diagnose the capability of MLLMs for video humor understanding. It contains a collection of funny videos collected from two complementary domains. Each clip is annotated with captions, descriptions, explanations, etc., supporting evaluation tasks such as caption matching and humor explanation. To broaden the applicability of v-Hub, we further construct an open-ended task, contributing to a comprehensive evaluation of MMLMs for video understanding. We evaluated a diverse range of MMLMs, spanning open-sourced and proprietary domains and covering specialized video LLMs and versatile OmniLLMs. Our findings reveal that current MLLMs heavily rely on linguistic cues for humor understanding, but are weak in deriving nuanced visual cues for understanding sophisticated video humor. Moreover, we empirically find that including audio is helpful for humor understanding, highlighting the informativeness of sound and the promise of incorporating rich modalities for complex video reasoning tasks.
Ethics statement
Our work follows widely recognized ethical principles in computing research, including the ACM Code of Ethics. In developing our benchmark, we considered the following aspects:
  • Contribute to Society and to Human Well-being: Our dataset is intended to advance multimodal AI research on humor understanding, a capability that has broad applications in safe content moderation, assistive technologies and cross-cultural communication. We employed three Human Intelligence Tasks (HITs) to gather data, and we fully considered cultural and linguistic diversity in humor to minimize potential negative impacts as much as possible, such as reinforcing harmful stereotypes, exposing sensitive content, or infringing on personal safety and privacy. To ensure accessibility and inclusivity, the dataset will be made broadly available for non-commercial research purposes.
  • Uphold High Standards of Scientific Excellence: We consistently adhere to principles of transparency and rigor throughout dataset curation and analysis. We meticulously document all preprocessing, annotation and evaluation procedures, and we will publicly release the relevant code to ensure independent verification and reproducibility. Furthermore, we provided fair compensation to all annotation personnel, ensuring their hourly wages exceeded the local minimum wage. Finally, we did not fabricate, falsify, or misrepresent data. Beyond the data annotators, no other human subjects were directly involved, and no personally identifiable information was used. Therefore, no additional ethics approval is required.
  • Avoid Harm: We carefully considered potential risks that could arise from constructing and releasing a humor video benchmark. To mitigate negative consequences, we implemented a multi-stage screening mechanism to exclude humorous content featuring violence, discrimination, or relying on stereotypes targeting vulnerable groups. Simultaneously, we adopted a three-person collaborative annotation scheme, ensuring that each data entry underwent three rounds of annotation. We explicitly document the limitations of the dataset to minimize unintended harm arising from cross-cultural or linguistic misunderstandings. Furthermore, we analyzed potential downstream risks, such as misuse for generating harmful or offensive humor, and explicitly warn against such applications in the dataset license.
  • Be Honest, Trustworthy and Transparent: We transparently disclose the characteristics, strengths, and limitations of the dataset. While this dataset thoroughly considers cultural and linguistic diversity in humor and encompasses diverse humor scenarios, it cannot cover all dimensions of global humor cultures. Furthermore, Charlie Chaplin’s Silent Films constitute a significant portion of the dataset, inevitably introducing potential cultural bias. We confirm that there are no conflicts of interest that may compromise the independence of our research, and all funding sources are clearly acknowledged. At the same time, we guarantee that we do not misrepresent related work, nor do we claim capabilities beyond what our benchmark enables.
  • Be Fair and Take Action not to Discriminate: We strived for fairness in both dataset curation and evaluation. And we made efforts to avoid humor that demeans particular groups. We also emphasize that the dataset should not be used to develop systems that discriminate, disenfranchise, or oppress individuals.
  • Respect the Work Required to Produce New Ideas and Artefacts: We credit the creators of ideas, inventions, work, and artefacts, and respect copyrights and property. All videos are sourced from publicly available materials with appropriate licenses. Where possible, we provide attribution to content creators and respect cultural heritage by excluding sensitive or protected media. Where possible, we will provide attribution to content creators and respect their work by excluding sensitive or protected media content.
  • Respect Privacy: Our work did not use private or personally identifiable data. All videos have been anonymized or utilize publicly available Charlie Chaplin silent films, carefully mitigating the risk of re-identification. Furthermore, the dataset is restricted to legitimate academic research under the dataset license.
  • Honour Confidentiality: Our work did not involve any confidential or proprietary information. Reviewers and collaborators were only provided with materials approved for release. We commit to maintaining confidentiality in peer review and in handling sensitive communications.
Appendix A Additional related work
A.1 From LLMs to Video LLMs
Large language models have demonstrated outstanding capabilities in many domains, including natural language processing, coding, math, and reasoning, ushering in new breakthroughs for video understanding technology. Video LLMs integrate visual encoders with LLMs, leading to a unified model to reason across video and language in the same language space (Wang et al., 2024a; Liu et al., 2024; Lin et al., 2023). Early video LLMs employ pre-trained image encoder and video encoder to encode only video frames  (Zhang et al., 2023; Maaz et al., 2023; Li et al., 2023; Lin et al., 2023). Recent works augment video LLMs with an audio encoder to align visual, auditory, and textual modalities in the same language space. Moreover, the audio encoder is supposed to capture diverse environmental sound apart from human speech since sound has been shown to contain amounts of commonsense knowledge  (Cheng et al., 2024; Xu et al., 2025).
Appendix B Crowdworking Details
B.1 Processing and filtering
Harmful content detection.
Before the annotation process began, we manually filtered out videos that contained potentially harmful content to ensure the video data’s safety and quality (Figure 4 visualizes our annotation interface). Based on the criteria outlined by Thoppilan et al. (2022), we defined 6 categories of harmful contents, following aspects are checked for each video.
  • Discrimination. Videos that display discrimination based on race, gender, sexual orientation, age, disability, appearance (e.g., obesity), or religion.
  • Animal Cruelty. Videos that depict the abuse or mistreatment of animals.
  • Dangerous Activities. Videos that include dangerous content such as drug use, criminal behavior, bullying, terrorism, rumor propagation, incitement, or misinformation.
  • Physical Violence. Videos containing acts of physical violence against individuals, including fighting, severe injuries, bleeding, self-harm, or torture.
  • Obscenities. Videos that contain explicit language, sexual behavior, or suggestive content.
  • Shocking Content. Videos that include startling or fear-inducing elements such as gunshots, explosions, or jump scares.
In addition to harmful content detection, videos are also evaluated based on their quality:
  • Confusing: Videos that are incomplete or otherwise difficult to understand.
  • Low Resolution: Videos with a level of clarity that makes it challenging to discern the content.
Chaplin video segmentation.
We selected 62 silent films by Charlie Chaplin and hired annotators to meticulously review each film, manually recording humorous moments to ensure each mime clip illustrates a whole mime through a single event or multi events. And we removed videos where both the reason for the humor and the action were repetitive (e.g. humor arising from a comical action due to inflexibility, such as failing to position a ladder properly) to ensure the quality and consistency of the videos and their annotations.
Speech reliance minimization.
To ensure reliable identification of humorous content, we instructed two annotators to independently review each video and confirm the presence of clear humor. Each annotator was also instructed to review each video and label whether humor was primarily conveyed through visual cues and could be understood independently of speech. Only videos for which both annotators agreed were retained for the final dataset. We further employed Whisper (Radford et al., 2023), a performant speech-to-text model, to transcribe audio. Since Whisper transcribes filler sounds (e.g., “uh,” “hmm”) and other minimal utterances, we excluded any videos where the transcribed text exceeded 10 characters. Additionally, videos containing non-English speech were retained but muted, removing dependence on linguistic cues.
B.2 Annotation
Annotator training. We provided appropriate annotation training for crowdworkers, offering detailed explanations of the annotation platform’s usage and the annotation guidelines for different tasks. Additionally, we supplied an annotation manual (Figure 6) and corresponding instructional videos, which included specific descriptions and examples of the annotation requirements for crowdworkers to consult at any time during the annotation process.
Qualification.The recruited crowdworkers were mainly from China, all possessing at least an undergraduate education and with English background. Before formal annotation began, we conducted training sessions and a qualification review. During the qualification stage, crowdworkers were required to annotate 15 video samples. We manually reviewed their results and assigned scores based on the annotation guidelines. Ultimately, we selected eight qualified annotators.
For the annotation process, we adopted a three-person collaborative annotation scheme, ensuring that each data entry underwent three rounds of annotation. First, an annotator performed the initial annotation. Next, a second annotator reviewed and supplemented the annotation. Finally, a third annotator reviewed and further refined the previous two rounds of annotations. The annotators rotated through these three roles, and each annotation round was tracked to ensure that the three rounds for each data entry were completed by different annotators. For humor rating and video captions, annotators were required to independently provide their own answers. For the remaining annotation tasks, when the second and third annotators reviewed and modified the previous annotations, they were required to submit a new annotation if they identified any issues. If a specific annotation issue was modified in all three rounds for a given video, we conducted a final review to assess the validity of the annotation results.
Figure 4: Interface for the Harmful Content Detection HIT.
Figure 5: Interface for HIT.
Figure 6: Interface for Annotation Manual for data annotation.
Appendix C Data Statistics
(a) Distribution of video lengths
(b) Distribution of humor types
Figure 7: Data statistics: video length and humor type distributions.
Duration.
All videos in our final dataset are restricted to a duration of 5–60 seconds, with the majority concentrated within 30 seconds (see Figure 7(a)). This design ensures that humor is self-contained, sufficiently nuanced, and compatible with the context length limits of most MLLMs.
Diversity.
To show our dataset contains a variety of humor, we follow Buijzen & Valkenburg (2004) to categorize humor into five categories (see Figure 7(b)): Slapstick humor, Clownish humor, Surprise, Irony, Misunderstanding and Others (e.g. Parody, Miscellaneous, Satire).
Visual-centric.
Our dataset is predominantly visual centric, with 99% of videos relying primarily on visual cues. Specifically, per the four groups defined in Section 2.4: Visual, Visual+Text, Visual+Audio, and Visual+Autio+Text, 600 videos (63%) fall into the category Visual, meaning humor is entirely derived from facial expressions, object interactions, or visual effects without reliance on text or sound effects. 214 videos (22%) integrate audio, indicating that while humor remains visually driven, auditory elements such as background music or sound effects enhance the comedic impact. 92 videos (9%) are classified as Visual+Text indicating that video humor is extended through additional visual text. 46 videos (5%) combines the three modalities to deliver humor. The remaining 1% includes videos, in which speech plays a minor role but does not dominate the expression of humor.
Figure 8: Video distribution over four groups.
Appendix D Additional Experimental Notes
D.1 DETAILS OF GENERATE OPEN-ENDED QA PAIRS
We employed GPT-4o to generate QA pairs for each video, with the questions primarily covering temporal, descriptive, and causal aspects. The specific prompts used for QA generation are provided in Table 16. Subsequently, annotators manually reviewed and revised the QA pairs for each video to ensure their accuracy and quality.
D.2 Details of evaluation methods
BERTScore.
It evaluates the semantic similarity between generated answers and ground truth (GT) references by leveraging contextual embeddings from pre-trained language models (Zhang* et al., 2020). Instead of relying on exact lexical overlap, BERTScore computes similarity at the token level in embedding space and aggregates these scores into three metrics:
  • Recall measures how well the GT tokens are semantically captured by the generated answer.
  • Precision measures how relevant the generated tokens are with respect to the GT reference.
  • F1 Score is the harmonic mean of precision and recall, providing a balanced assessment of semantic alignment.
The inclusion of BERTScore allows us to evaluate whether model-generated answers semantically align with human references, thereby assessing the adequacy and quality of humor comprehension.
Human evaluation.
We randomly sampled 50 explanations generated by models for human evaluation. To ensure consistency in the evaluation criteria, we assigned one annotator to rate the humor explanations generated by models. The scores ranged from 0 to 100, and were subsequently normalized by a factor of 100, yielding final results within the range $[0,1]$.
D.3 Baseline Models
Table 6: Evaluated Models.
Models #Parameter Proprietary Input Modality
Text Video Video+Audio
Qwen2.5-VL 72B [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Qwen-2.5-Omni 7B [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Intern3.5-VL 8B [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MiniCPM2.6-o 8B [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Video-SALMONN-2 7B [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GPT-4o - [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Gemini-2.5-flash - [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
To evaluate multimodal large language models’ ability to understand video humor, we selected state-of-the-art models representing three distinct input modalities, as summarized in Table 6. Specifically, we include multimodal LLMs that process raw visual frames and text, and omni LLMs that integrate both text, video and audio signals. This set covers both public models (e.g., Qwen2.5-VL-72B, Intern3.5-VL) and proprietary models (e.g., Gemini-2.5-flash, GPT-4o), offering a broad perspective on current approaches. Each model is evaluated under all input conditions it can handle (see Section 4.1): for instance, omni-modal models can participate in the Text-Only, Video-Only, and Video+Audio groups, whereas multimodal models are tested exclusively with textual input and raw visual frames. This setup allows us to isolate how each model category—multimodal and omni-modal—contributes to humor understanding across diverse input modalities.
Appendix E Additional Experimental Results
Table 7: Comparison between MLLMs and their base LLMs under the Text-Only setting.
Models Open-ended QA
SentBERT METEOR Precision Recall F1 Score
Qwen2.5-VL-72B 0.792 0.622 0.744 0.740 0.738
Qwen2.5-72B 0.710 0.636 0.680 0.723 0.700
Qwen2.5-Omni-7B 0.740 0.555 0.703 0.703 0.698
Qwen2.5-7B 0.692 0.539 0.657 0.688 0.667
MLLMs vs. their base LLMs.
MLLMs are usually derived from a pre-trained base LLM by adding a visual encoder or multimodal modules. For instance, Qwen2.5-VL-72B extends Qwen2.5-72B, and Qwen2.5-Omni extends Qwen2.5-7B (see Table 7). In the Text-Only setup, Qwen2.5-Omni surpasses Qwen2.5-7B with a SentBERT score of 0.740 (vs. 0.692) and an F1 score of 0.698 (vs. 0.667) on Open-ended QA task, suggesting that multimodal training can confer advantages even when only textual descriptions are available—possibly because the model has learned richer contextual associations during training. Please refer to Table 11 for more details on humor explanation and caption matching tasks.
Table 8: The impact of requiring background knowledge support on video humor understanding.
Models Explanation Matching Open-ended QA
SentBERT METEOR F1 Score Accuracy SentBERT METEOR F1 Score
Gemini-2.5-flash 0.503 0.211 0.568 0.633 0.437 0.268 0.550
video-SALMONN-2 0.273 0.153 0.513 0.266 0.306 0.159 0.520
MiniCPM2.6-o 0.404 0.166 0.532 0.378 0.318 0.103 0.442
Qwen-2.5-Omni 0.404 0.157 0.531 0.571 0.384 0.105 0.486
Background knowledge does not necessarily improve video humor understanding.
The results in Table 2 and Table 8 show that there is no significant difference between the mean scores and the scores for videos that need background knowledge to perceive humor under Video+Audio setting. For example, under the Video+Audio setting, the Gemini-2.5-flash attains an average F1 score of on the Explanation task for Background-Dependent videos 0.568, which is statistically similar to its F1 score of 0.550 on the full dataset. This suggests that the language-model component of MLLMs already encodes most of the cultural background knowledge necessary for humor comprehension, meaning that the absence of explicit background knowledge in the input does not significantly degrade their performance. MLLMs do not show a significant disadvantage in understanding videos that require background knowledge compared to those that do not, potentially because video humor rarely relies on specific background knowledge, making it universally understandable.
Proprietary MLLMs show stronger resilience to multimodal inputs compared to public MLLMs.
The results in Table 9 indicate that current MLLMs rely heavily on linguistic cues to generate reasonable explanations, and struggle to effectively extract semantic information from raw visual or auditory signals. For example, Qwen2.5-VL-72B attains a preference score of 0.687 under Text-Only, significantly outperforming its Video-Only score of 0.423. Furthermore, although closed-source models demonstrate greater robustness under multimodal inputs, they still struggle to align visual and audio cues to enhance humor comprehension. For instance, Gemini-2.5-flash achieves 0.546 (Video-Only) and 0.566 (Video+Sound).
Appendix F The Use of Large Language Models
In this work, we used LLMs as assistive tools in several stages of the research:
  • Dataset construction. We initially employed GPT-4o to assist in generating candidate QA pairs and humor categories from video content. All outputs were subsequently reviewed and revised by human annotators to ensure correctness and quality.
  • Code assistance. LLMs were used to help generate parts of the evaluation code, which were then verified and refined by the authors.
  • Writing support. LLMs were used to write and polish some sentences in the paper for readability.
Table 9: Human preference comparison of humor explanations across four model categories.
Models Proprietary Type Setting
Text-Only Video-Only Video+Audio
Qwen2.5-VL-72B [Uncaptioned image] MLLM 0.687 0.423
Qwen2.5-Omni [Uncaptioned image] OmniLLM 0.574 0.430 0.381
GPT-4o [Uncaptioned image] MLLM 0.654 0.576
Gemini-2.5-flash [Uncaptioned image] OmniLLM 0.651 0.546 0.566
Table 10: Model performance on Humor Explanation, Caption Matching, and Open-ended QA.
Explanation Matching Open-ended QA
Models Precision Recall F1 Score Accuracy Precision Recall F1 Score
Text-Only
Gemini-2.5-flash 0.550 0.616 0.580 0.611 0.710 0.744 0.723
video-SALMONN-2 0.570 0.612 0.589 0.367 0.650 0.636 0.639
MiniCPM2.6-o 0.519 0.614 0.562 0.518 0.460 0.666 0.543
Qwen-2.5-Omini 0.540 0.605 0.570 0.644 0.703 0.703 0.698
Qwen-2.5-VL-72B 0.548 0.615 0.578 0.719 0.744 0.740 0.738
Intern3.5-VL 0.548 0.618 0.580 0.632 0.688 0.701 0.689
GPT-4o 0.546 0.626 0.581 0.767 0.698 0.746 0.718
Video-Only
Gemini-2.5-flash 0.522 0.583 0.550 0.583 0.547 0.570 0.556
video-SALMONN-2 0.487 0.525 0.504 0.259 0.549 0.509 0.525
MiniCPM2.6-o 0.500 0.543 0.520 0.364 0.397 0.522 0.452
Qwen-2.5-Omini 0.503 0.543 0.521 0.55 0.538 0.457 0.488
Qwen-2.5-VL-72B 0.526 0.571 0.547 0.673 0.567 0.541 0.550
Intern3.5-VL 0.522 0.568 0.543 0.640 0.544 0.545 0.542
GPT-4o 0.515 0.585 0.547 0.667 0.538 0.579 0.556
Video+Sound
Gemini-2.5-flash 0.521 0.583 0.550 0.588 0.543 0.570 0.554
video-SALMONN-2 0.487 0.529 0.506 0.255 0.556 0.526 0.538
MiniCPM2.6-o 0.501 0.553 0.523 0.404 0.494 0.541 0.513
Qwen-2.5-Omini 0.509 0.558 0.531 0.623 0.565 0.507 0.529
Table 11: Comparison between MLLMs and their base LLMs under the Text-Only setting.
Models Explanation Matching
SentBERT METEOR Precision Recall F1 Score Accuracy
Qwen2.5-VL-72B 0.553 0.250 0.548 0.615 0.553 0.719
Qwen2.5-72B 0.546 0.245 0.552 0.615 0.581 0.646
Qwen2.5-Omni-7B 0.547 0.232 0.605 0.570 0.547 0.644
Qwen2.5-7B 0.568 0.241 0.540 0.613 0.573 0.522
Table 12: The impact of requiring background knowledge support on video humor understanding.
Models Explanation Matching Open-ended QA
Precision Recall F1 Score Accuracy Precision Recall F1 Score
Gemini-2.5-flash 0.545 0.595 0.568 0.633 0.540 0.566 0.550
Qwen-2.5-Omni 0.517 0.547 0.531 0.571 0.533 0.460 0.486
MiniCPM2.6-o 0.516 0.550 0.532 0.378 0.396 0.521 0.442
video-SALMONN 2 0.501 0.527 0.513 0.266 0.542 0.507 0.520
Table 13: The impact of background knowledge on video humor understanding.
Explanation Matching Open-ended QA
Models Precision Recall F1 Score Accuracy Precision Recall F1 Score
w/ Background Knowledge
video-SALMONN-2 0.561 0.568 0.563 0.331 0.554 0.525 0.535
MiniCPM-2.6-o 0.533 0.585 0.557 0.445 0.508 0.553 0.520
Qwen-2.5-Omni 0.536 0.582 0.557 0.667 0.583 0.540 0.556
w/o Background Knowledge
video-SALMONN-2 0.499 0.533 0.515 0.261 0.545 0.517 0.528
MiniCPM-2.6-o 0.521 0.563 0.540 0.412 0.490 0.537 0.509
Qwen-2.5-Omni 0.462 0.181 0.545 0.630 0.445 0.158 0.526
Table 14: The impact of video era on video humor understanding.
Explanation Matching Open-ended QA
Models Precision Recall F1 Score Accuracy Precision Recall F1 Score
Charlie Chaplin’s Silent Films
Gemini-2.5-flash 0.517 0.570 0.541 0.562 0.529 0.567 0.545
video-SALMONN-2 0.492 0.528 0.509 0.165 0.521 0.511 0.513
MiniCPM2.6-o 0.492 0.527 0.508 0.307 0.431 0.535 0.470
Qwen-2.5-Omni 0.495 0.528 0.510 0.494 0.515 0.482 0.493
User-Generated Funny Video
Gemini-2.5-flash 0.525 0.588 0.553 0.592 0.553 0.572 0.560
video-SALMONN-2 0.485 0.523 0.503 0.296 0.559 0.509 0.529
MiniCPM2.6-o 0.503 0.549 0.524 0.385 0.405 0.518 0.446
Qwen-2.5-Omni 0.506 0.549 0.525 0.571 0.547 0.448 0.487
Appendix G Prompts
We list our prompt in Tables˜15, 16, 19, 20, 22, 21, 18, 17, 25, 24 and 23
Table 15: Prompt for writing captions of videos.
And I will provide a description of the video and a list of descriptive captions that break down what happens in it.
Your task is to write a caption in one sentences from the video creator’s perspective — something you would write to attract viwers.
Requirements:
Please ensure it is related to the video content.
- Write as if you’re sharing it with an audience (e.g., use ’this’ or ’me’ naturally).
Output format:
Caption: <caption>
Video description: {video_description}
Descriptive captions: {descriptive_captions}
Table 16: Prompt for generate QA pairs.
These are frames from a video.
And you’ll be given a description of a video and an explanation of why it’s humorous to watch.
[…]
Table 17: Prompt for video QA.
System: You are a helpful AI assistant. You can analyze videos and answer questions about their content. Respond with short and concise answers. Avoid using unpronouncable punctuation or emojis.
User: These are frames from a video. Based on these frames, answer the following ${question} \n\n$
[…]
Table 18: Prompt for video explanation.
System: You are a helpful AI assistant specialized in video understanding and humor analysis. You can explain jokes clearly and naturally based on video content and video description. Please respond with short and concise answers. Avoid using unpronouncable punctuation or emojis.
[…]
Table 19: Prompt for video caption matching.
System: You are a helpful AI assistant. You can analyze videos and answer questions about their content. Please only output in the specified format. No extra text.
[…]
Table 20: Prompt for video with description QA.
System: You are a helpful AI assistant. You can analyze videos and answer questions about their content. Respond with short and concise answers. Avoid using unpronouncable punctuation or emojis.
User: You’ll be given a description of the video. Based on this information, answer the following ${question} \n\n$
[…]
Video Description: {video_description}
Table 21: Prompt for video with description explanation.
System: You are a helpful AI assistant specialized in video understanding and humor analysis. You can explain jokes clearly and naturally based on video content and video description. Please respond with short and concise answers. Avoid using unpronouncable punctuation or emojis.
[…]
Video Description: {video_description}
Table 22: Prompt for video with description caption matching.
System: You are a helpful AI assistant. You can analyze videos and answer questions about their content. Please only output in the specified format. No extra text.
[…]
Table 23: Prompt for video with sound QA.
System: You are a helpful AI assistant. You can analyze videos and answer questions about their content. Respond with short and concise answers. Avoid using unpronouncable punctuation or emojis.
User: Here’s a humorous video. Based on the its visual and audio information, answer the following ${question} \n\n$
[…]
Table 24: Video with sound explanation.
System: You are a helpful AI assistant specialized in video understanding and humor analysis. You can explain jokes clearly and naturally based on video content and video description. Please respond with short and concise answers. Avoid using unpronouncable punctuation or emojis.
[…]
Table 25: Video with sound caption matching
System: You are a helpful AI assistant. You can analyze videos and answer questions about their content. Please only output in the specified format. No extra text.
[…]